OpenStudy (praxer):

if the diagonals of a parallelogram are 3i^ + j^ -2K^ and i^ -3j^ +4K^ then how can we find the area.

OpenStudy (praxer):

@Callisto please help me..

OpenStudy (callisto):

Suppose we have to vectors a and b, which form the parallelogram. Since the given vectors are the diagonal of the parallelogram, a + b = <3, 1, -2> a - b = <1, -3, 4> Solve for a and b. With a and b, I think you can find the area of parallelogram using cross product.

OpenStudy (praxer):

@Loser66 can u hlp me with this please..

OpenStudy (loser66):

what grade are you? Do you know cross product?

OpenStudy (praxer):

i know cross product but i didn't get the explanation above... So, i am asking

OpenStudy (loser66):

ok, just follow him, and ....we find the meaning of the logic then, Ok?

OpenStudy (praxer):

ok.

OpenStudy (loser66):

get a, b from his equation first, what do you have?

OpenStudy (praxer):

ok a+b is<4,-2,2> a-b is <2,4,-6> then?

OpenStudy (praxer):

ok i get it he was saying like a+b to be the sides which add to give one diagonal and a-b to be the other is it so?

OpenStudy (loser66):

let a =(a1,a2,a3) b =(b1,b2,b3) a+b = (a1+b1, a2+b2,a3+b3) = (4,-2,2) a -b = (a1-b1,a2-b2,a3-b3)= (2,4,-6) so, you have a1+b1= 4 a1-b1 = 2 solve it, you have a1=3,b1= 1 got me so far?? do the same to get a2, b2, and a3 b3

OpenStudy (praxer):

is it like a+b = <3,1,-2> a-b =< 1,-3,4> now a+b+a-b = 2a = < 4, -2,2> now a=< 2,-1,1> b=<1,2,-3> now from cross product rest can be done. Is this correct???? @Loser66

OpenStudy (loser66):

wow, impressive, you got it!!! 

OpenStudy (praxer):

Sorry for the annoyance :). Anyways thank u :)

OpenStudy (praxer):

ohk i will

OpenStudy (loser66):

oh, you are right, my bad, I type the wrong numbers. hehehe.. sorry.

OpenStudy (loser66):

you are on the right track and you understand why you do so, right?

OpenStudy (praxer):

Yup, i understood it..

OpenStudy (loser66):

good luck

OpenStudy (praxer):

Thank you ! :)

OpenStudy (loser66):

np

OpenStudy (praxer):

I suppose u wont require a good luck, you are so good in mathematics... :)

OpenStudy (loser66):

nope, I always failed at the silly mistake which ruins all my stuff.

OpenStudy (praxer):

Talent matters. At least you can proceed. But i am sicked of maths.. Anyways i shall practice from 2day onwards.